OpenStudy (anonymous):

factor 9x^3 -54x^2+72x

OpenStudy (solomonzelman):

Take the 9 out,

OpenStudy (anonymous):

could you solve this step by step?

OpenStudy (solomonzelman):

9x^3 -54x^2+72x 9(x^3 -6x^2+8x) Hint: factors of 8 1 & 8 2 & 4 the constants will be both negative because the add up to negative number and multiply to a positive number. the constants have to add up -6 and multiply to 8 -2 + (-4)= -2-4=-6 and (-2) * (-4)= 8 so 9 (x-2) (x-4)